OpenStudy (anonymous):

Sean currently has an account balance of $3,981.04. He opened the account five years ago with a deposit of $3,760.19. If the interest compounds daily, what is the interest rate on the account?

OpenStudy (wolf1728):

See attached graphic log(3,981.04/3,760.19)/365*5 = 0.0247867511 / 1,825 = 1.35817814407889E-005 log(1 + [rate/n]) = 1.35817814407889E-005 10^1.35817814407889E-005 = log(1 + [rate/n]) 1.0000312737 = 1 + [rate/n] .0000312737 = (rate/n) Multiply by 365 0.0114148992 rate = 1.14148992% and a calculator to check that is here: http://1728.org/compint.htm